OpenStudy (anonymous):

The half-life of Palladium-100 is 4 days. After 16 days a sample of Palladium-100 has been reduced to a mass of 3 mg. What was the initial mass (in mg) of the sample? What is the mass 8 weeks after the start?

OpenStudy (anonymous):

y = initial mass * ( 1/2) ^ ( t / halflife)

OpenStudy (anonymous):

y = yo * ( 1/2) ^ ( t/4)

OpenStudy (anonymous):

3= yo (1/2)^(16/4)

OpenStudy (anonymous):

3/ ( .5^4) = initial mass

OpenStudy (anonymous):

y = 48 * ( 1/2) ^ ( t/4)

OpenStudy (l):

16/4 = 4 3 * 2^4 = 3 * 16 = 48mg -- initial mass 8 weeks after the start = 8*7 = 56 56/4 = 14 48*(0.5)^14 ~ 0,002929688 *

OpenStudy (anonymous):

thank you